OpenStudy (yanasidlinskiy):

So guys! I need some quick help and then I'm out! It's urgent and I can't resist not posting it here!

OpenStudy (yanasidlinskiy):

\[\frac{ 2 }{ 5 } (4x+6) = 2 (2-3x) + 6\]

OpenStudy (yanasidlinskiy):

My answer came out to something funky like 40x + 10 = 10 - 6x

OpenStudy (yanasidlinskiy):

Idk what I did wrong.

ganeshie8 (ganeshie8):

long time!

OpenStudy (yanasidlinskiy):

I know!!! Omg!! You're amazing!!

ganeshie8 (ganeshie8):

\[\frac{ 2 }{ 5 } (4x+6) = 2 (2-3x) + 6\] multiply \(5\) through out and get \[2 (4x+6) = 10 (2-3x) + 30\] distributing \[8x+12 = 20-30x + 30\]

OpenStudy (yanasidlinskiy):

Wait. So once I multiply I don't put it back as a fraction like: \[\frac{ 8 }{ 5 } x + \frac{ 12 }{ 5 }\]

ganeshie8 (ganeshie8):

we have \[\frac{ 2 }{ 5 } (4x+6) = 2 (2-3x) + 6\] multiplying \(5\) through out gives \[\color{red}{5}*\frac{ 2 }{ 5 } (4x+6) = \color{red}{5*}2 (2-3x) + \color{red}{5*}6\]

OpenStudy (phi):

**Wait. So once I multiply I don't put it back as a fraction like:*** you could do that, by gan is showing a nicer way.. it gets rid of the fraction

OpenStudy (phi):

the idea is multiply both sides by 5 (which is OK to do) and on the left, the 5/5 cancel. and no more fraction

OpenStudy (phi):

btw, you should continue to simplify until you get x= some number.

OpenStudy (yanasidlinskiy):

Yes, that's what I was doing. I got 1! Thanks guys for the help! I appreciate it! I'll use this as an example to lead me through more! Thanks again!

OpenStudy (yanasidlinskiy):

I wish you were my teacher, ganeshie! It's soo much easier with your steps!!

ganeshie8 (ganeshie8):

one key thing about math in general is - follow the rules and all your methods will produce the same answer in the end and ofcourse there always exists the best way to do any problem :)

OpenStudy (yanasidlinskiy):

Yes, that's what I was doing. I was taught some other way like finding LCD which drives me insane!!! I always want to die when I see finding LCD. It's a killer! Oh yes! That's why I always ask you, because you always have some way that I do not know!!;))